Find the particular solution of the differential equation 8d2x/dt2 + 16dx/dt -5x =0, when x(0) = 16 and x'(0) = -8.


1
Expert's answer
2021-08-12T18:15:50-0400

Given IVP:


"8x''(t)+16x'(t)-5x(t)=0"

"x(0)=16,\quad x'(0)=-8"

The characteristic equation is as follows

"8k^2+16k-5=0"

Roots:

"k_1=-1-\sqrt{26}/4,\quad k_1=-1+\sqrt{26}/4"

General solution of IVP:

"x(t)=Ae^{k_1t}+Be^{k_2t}=e^{-t}(Ae^{-\sqrt{26}/4t}+Be^{\sqrt{26}/4t})"

The initial conditions give

"x(0)=A+B=16\\
x'(0)=-(A+B)+\sqrt{26}/4(-A+B)=-8""A=-8/13(-13+\sqrt{26}), \quad B=8/13(13+\sqrt{26})"

Finally,

"x(t)=e^{-t}\Big(-8/13(-13+\sqrt{26})e^{-\sqrt{26}/4t}\\+8/13(13+\sqrt{26})e^{\sqrt{26}/4t}\Big)"